OpenStudy (anonymous):

What is the ratio of x to y? 4x = 5y @chaser71 Answers below

OpenStudy (anonymous):

To cross multiply and end up with 4x = 5y, it could have been x/y = 5/4 so the ratio of x to y = 5 to 4

OpenStudy (anonymous):